OpenStudy (anonymous):

Help!!!!!!! Divide. Assume that no denominator equals zero 8d^3/6c divided by c^3d^4/24

OpenStudy (faiqraees):

what is the exponent on c

OpenStudy (anonymous):

c^3

OpenStudy (faiqraees):

\[\large\rm \frac{8d^3}{6c}\div \frac{c^3d^4}{24}\] \[\large\rm \frac{8d^3}{6c}* \frac{24}{c^3d^4}\] \[\large\rm \frac{8d^3*24}{6c*c^3d^4}\] Can you work out the rest?

OpenStudy (anonymous):

no i didnt get it..

OpenStudy (anonymous):

is it 64c^2/d

OpenStudy (faiqraees):

Nope wrong. Which part in the given steps you didnt understand

OpenStudy (anonymous):

how bout 16c^2/d

OpenStudy (faiqraees):

I think you are not multipying by the reciprocal

OpenStudy (anonymous):

it must be 32/c^4d

OpenStudy (faiqraees):

correct

OpenStudy (anonymous):

ohh makes sense! can you help me with another?

OpenStudy (anonymous):

d^2-1/d^2-d divided by d-1/d

OpenStudy (anonymous):

assume that no denominator equals zero

OpenStudy (anonymous):

@Qwertty123

OpenStudy (qwertty123):

@FaiqRaees was helping you just fine! :D Can you help again?

OpenStudy (anonymous):

i need help with the question right before i tagged you.. its different

OpenStudy (anonymous):

i need help with the question right before i tagged you.. its different